# Solar photospheric velocities measured in space: a comparison between SO/PHI-HRT and SDO/HMI

### v1

@@
The Polarimetric and Helioseismic Imager (SO/PHI) onboard Solar Orbiter is a spectropolarimeter scanning the Fe i line at 617.3 nm, providing data of the solar photosphere. The same line is sampled by the Helioseismic and Magnetic Imager (HMI) on board the Solar Dynamics Observatory (SDO) and many other on-ground instruments. In this paper, we aim at assessing the consistency between line-of-sight (LoS) velocity measurements from the two instruments. Reliable measurements of up and down flows from SO/PHI are crucial and unique when Solar Orbiter is facing the far side of the Sun. Also, a combination of measurements from two vantage points to study horizontal flows must rely on consistent observations. For this purpose, we compare the LoS velocity measured by SO/PHI’s High Resolution Telescope (SO/PHI-HRT) and SDO/HMI on 29 March 2023, when Solar Orbiter was crossing the Sun-Earth line at a distance of 0.39 au from the Sun. Because such co-alignments are rare, this configuration offered an almost unique opportunity to directly compare data products from both telescopes. The data are aligned and remapped to allow a pixel-by-pixel comparison of the whole time series of 4 hours length. Temporal and spatial variations are considered for a direct combination of the measurements. The LoS velocity distributions are evaluated and a clear linear relation is found between the two instruments with a slope of 0.94 and a correlation of 90%. We find that the signals form at similar heights, with a separation of 1012 km, which is larger than previous estimates. A close-up look at the penumbra of a sunspot and its Evershed flow is presented. We conclude that the signals inferred by SO/PHI-HRT and SDO/HMI show very good agreement and high correlation when instrumental effects and large-scale velocities on the Sun are properly accounted for.
@@
Despite the clear advantages of satellite observations, such as the absence of the disturbance from the Earth’s atmosphere and the possibility of uninterrupted data acquisition, many challenges have to be faced to achieve a consistent data reduction. In the case of SO/PHI the highly eccentric orbit and low telemetry periods when the spacecraft is at more than 1 au from Earth, make the SO/PHI data reduction considerably more challenging. On the other hand, SO/PHI can for the very first time measure the photospheric properties of the Sun from different viewpoints than Earth-bound observatories (Müller et al. 2020). This unique orbit opens the possibility for novel scientific applications, such as building fast synoptic maps of the LoS magnetic field (Loeschl et al. 2024b, a) or direct combination of intensity and magnetic field data in stereoscopic configuration (e.g. Valori et al. 2023; Albert et al. 2023; Romero Avila et al. 2024). Comparisons of the magnetic field vector between SO/PHI and SDO/HMI have been carried out by Sinjan et al. (2023) and Moreno Vacas et al. (2024) to assess the similarity between the data products of these two instruments. Such inter-instrument comparisons are fundamental for combining data from different vantage points, particularly when tracking features on the far side of the Sun (e.g. Finley et al. 2025).
@@ 2 Data
The data products chosen for this comparison are the 45 s cadence data products of SDO/HMI (based on the so-called MDI-like algorithm, Couvidat et al. 2016) and the newly improved standard results of the SO/PHI-HRT pipeline (Sinjan et al. 2022; Kahil et al. 2022; Bailén et al. 2024), which is now also correcting for residual stray light, based on the inversion of the radiative transfer equations (RTE) in Milne-Eddington approximation with the MILOS code (Orozco Suárez & Del Toro Iniesta 2007). This data product became the standard data product in most recent data releases111Fourth SO/PHI-HRT data release notes. (see Sinjan et al. in preparation, for more details).
@@ 4.1 Fourier space
The power spectra of the remapped SO/PHI-HRT and SDO/HMI time series and their phase difference and coherence spectrum are shown in Fig. 2. The phase difference has been corrected for the time difference between the two time series. The coherence spectrum shows the linear relation between two signals and it is not affected by their amplitude (Vigeesh et al. 2017). The first thing to notice is the similarity between the power spectra of the two instruments, which is made more evident by the diagrams showing the phase difference and the coherence between the two time series. The phase difference remains consistently close to 0∘ across most of the domain, while the coherence spectrum shows very high values (above 80%) along the p-modes and f-mode ridges and in the internal gravity waves and granulation regime. The phase difference diagrams can be used to estimate the average height difference between the two signals (see equations (2) and (3) in Vesa et al. 2023). We select the area of the diagram where the waves have real vertical wavenumber, which means that they propagate and they are not evanescent (as when the vertical wave number is imaginary). We can then restrict our choice only to vertically propagating waves by considering only small horizontal wavenumbers (see the red dashed box in the bottom right panel in Fig. 2). We assume the cut-off frequency, buoyancy frequency, and sound speed to be equal to 5.4 mHz, 4.9 mHz, and 7 km/s respectively. We find that the SO/PHI-HRT signal forms approximately 10 km above SDO/HMI, with a broad distribution between -10 and 30 km. Previous result by Schou et al. (2023) suggested a formation height of ~156 km for SDO/HMI and ~158 km above the solar surface for SO/PHI obtained with the help of a MURaM simulation. This result was obtained by retrieving the velocity through the MDI-like algorithm also for SO/PHI, so we would expect those signals to be closer than those addressed in that work.
@@ 4.2 Distributions
As a second assessment of the similarity of the SDO/HMI and SO/PHI-HRT Dopplergrams, a scatter plot of the remapped SO/PHI-HRT and the non-interpolated SDO/HMI time series is shown in Fig. 3. Only the datasets with time difference of 5 s are selected for plotting in this figure because the width of the distribution is affected by the time discrepancy. The colors correspond to the number of points inside a specific bin in logarithmic scale. Two functions have been fit to this distribution: a linear function and a 2-D Gaussian distribution. The linear fit shows a clear correspondence between the data, with a slope of 0.94, a correlation coefficient of 90%, and an offset of -203 m/s. The Gaussian distribution indicates the spread between the two signals, which is 132 m/s at 1. We also determine the mean of the distribution, which provides information on the offset of the zero point between both time series. The average of the distribution for SDO/HMI is at -195 m/s, whereas the SO/PHI-HRT mean value is at -7 m/s. Discussing the absolute calibration of these instruments is a challenge that goes beyond the scope of this work, but it is indeed a point that needs to be mentioned. In the SDO/HMI case, the reference velocity can be set in many ways available in the literature (e.g. Mathew et al. 2003; Grigor’ev et al. 2011; Okamoto et al. 2009; de la Cruz Rodríguez et al. 2011; Welsch et al. 2013), including simply setting the mean signal at the center of the Sun to zero (after the removal of all the LSC). The same approach is not always possible for SO/PHI-HRT, due to its smaller FoV and the fact that it often points such that solar disk center does not lie in its FoV.
@@
To investigate the presence of any temporal dependence, we repeated the analysis described above for each of the 240 frames separately. Fig. 4 shows a selection of fit parameters as a function of time. The top panel displays the results of the linear fit on each of the 240 observation pairs of frames. The blue points show the slope and the orange points the offset obtained by the linear fit. We can clearly distinguish a linear trend through the 4 hours of observation in which both values evolve towards a better concordance. The middle panel depicts the coordinates of the average of the Gaussian distribution. The left axis (in green, coordinates of the center of the distribution) corresponds to the mean values of each dataset of the remapped SO/PHI-HRT, the right coordinates (in red, coordinates of the center of the distribution) to SDO/HMI. The range of the two -axes is the same for both coordinates, meaning that the relative variation with time of both values is directly comparable. In this case, the SDO/HMI values do not show any significant variation with time, while the remapped SO/PHI-HRT values go from 0 to -15 m/s. The bottom panel shows the root mean square (rms) of the remapped SO/PHI-HRT (purple) and SDO/HMI (brown) time series, which is used to quantify the amplitudes measured by the instruments. The two axes in this case are exactly the same for both distributions. The remapped SO/PHI-HRT rms is overall higher than SDO/HMI as expected by the slope of the linear fit shown in Fig. 3 being smaller than one. On the other hand, we see a decrease of the rms in SO/PHI-HRT of ~5 m/s and an increase in SDO/HMI of ~10 m/s, which is also consistent with the increase of the slope in time.
@@
In a further step we look at the spatial variation of the results of the linear fit. For that, the FoV is divided into square regions of 88 pixels and we perform a linear fit in each of these regions through the whole time series. We also compute the averaged correlation between the two time series in each sub-region. The results are shown in Fig. 5. The top panel shows the spatial-distribution of the correlation, which is overall very high with a mean value of 91%. Low values are found especially within the sunspot umbra, where the correlation drops to 60-70%. Many factors lead to this low correlation including the low intensity signal, the large line splitting due to the strong magnetic field that can become as large as the total wavelength range in the line that is scanned, the small velocity in the umbra, the few measured wavelength points within the line (which are different for SDO/HMI and SO/PHI), different algorithms to retrieve the velocity signal, all contribute to the low correlations between the two instruments within umbrae (see Sect. 5).
@@ 4.3 Sunspot penumbra
In this section the penumbra of the sunspot is analyzed in more detail. The penumbra is defined as those pixels with intensity higher than 55% and lower than 89% of the averaged intensity continuum in a nearby quiet Sun area. A reproduction of the scatter plot between the remapped SO/PHI-HRT and the SDO/HMI time series, focusing exclusively on the penumbra and on its Evershed flow (Evershed 1909) is visible in Fig. 6. The contours defining the penumbra boundaries are also visible in the lower right part of the same figure. It is evident that the scatter of the points is smaller than in Fig. 3, which can be explained by the smaller FoV considered here (and thus smaller spatial variation) and is well represented by the higher correlation coefficient (95%) with respect to that of the full FoV. Another remarkable result is the slope of the linear fit, which does not differ from the one obtained by fitting the data of the whole FoV. This is unlikely to be by chance given the small error bars associated with each data point. Additionally, the offset between the two signals is smaller than that found in Fig. 3. This is expected from the position of the sunspot in the bottom panel of Fig. 5, where it lies in an area where the offset is smaller than the average (as obtained by the linear fit shown in Fig. 3).
@@
The middle panel of Fig. 7 displays the velocity on the elliptical polar grid. Note that the velocities plotted in this panel are temporal averages over the full 240 frames recorded by the SO/PHI-HRT. The two branches of the Evershed flow are clearly visible in the left and right part of the plot, which consists of the limb (outlined by the red box) and disk center (blue box) oriented sides of the penumbra, respectively. In order to compare the values obtained by the two instruments, we perform an azimuthal average (along the axis) in the red and blue boxes for each time series. The results are shown in the right panel of Fig. 7, where the axis represents the radial distance on the elliptical polar grid. The colors are the same as of the boxes shown in the middle panel, the crosses show the result obtained with the remapped SO/PHI-HRT and the circles those from SDO/HMI. The curves resulting from the data of the two instruments lie very close to each other until the penumbra-quiet Sun boundary (which lies roughly at a distance of 15 px along the axis). The asymmetry between the two sides of the penumbra (limb- and disk center-side, e.g. Schlichenmaier et al. 2004) is to be expected. Also, a difference in the flows can be caused by the asymmetric shape of the sunspot.
@@ 5 Discussion
We have reported a comparison of the LoS velocity measured by SO/PHI-HRT and SDO/HMI while Solar Orbiter was in inferior conjunction. The data acquired by SO/PHI-HRT have been post-processed and remapped onto the SDO/HMI detector frame to allow a pixel-by-pixel comparison. Despite differences in instrumentation, calibration, cadence, and velocity retrieval methods, the comparison shows an excellent agreement of the LoS velocity signals obtained with these two telescopes. We measure a height difference of 1012 km between the two time series and obtain a slope of 0.94 for the linear fit of the scatter plot between the two instruments and a correlation coefficient of above 90%. The measured height difference is larger than the result obtained from a MURaM simulation by Schou et al. (2023) and it is possible to estimate the velocity variation due to such a separation. If we consider the conservation of mass from one layer to the other, it holds
where is the density at each layer and is the vertical velocity. A ratio of 0.95 for 10 km height separation is obtained by considering the atmospheric densities from the VAL-C model (Vernazza et al. 1981), whereas 11 km of separation would be necessary to obtain a ratio equals to 0.94. Considering the broad distribution of the height separation between the two instruments, it is not possible to address it as the unique cause for having a ratio smaller than 1. One more reason is to be found in the bottom panel of Fig. 4, which shows that the rms of the remapped SO/PHI-HRT is always higher than that of SDO/HMI. This is due to the spatial deconvolution of the SO/PHI-HRT PSF including the stray light component during the data reduction, a correction that is not included in the SDO/HMI pipeline, but it can be made available on request. We compare these data with the velocities from SDO/HMI without stray light removed, as we want to compare the standard data products, which are the most likely to be used for studies comparing or combing velocities from the two instruments. One more element that could also contribute is the different sampling of the spectral line. SO/PHI-HRT samples the line in five points with a spectral resolution of 58000, SDO/HMI in six points with a spectral resolution of 81000. Stief et al. (2019) showed that the velocity retrieved with different sampling and at different points along the line can substantially affect the measurement. However, a lower spectral resolution reduces the variation in the retrieved velocity near the core of the line (see the yellow curve in Fig. 5 in Stief et al. 2019 for the SDO/HMI case), meaning that we do no expect this to be the main component of the difference between the two instruments.
@@
The spatial correlation and dependence of the linear fitting results are shown in Fig. 5. The correlation map shows very high values through the whole FoV with an average of 91%, with the exception of the sunspot umbra, where the measured LoS velocities are generally lower than in the penumbra or quiet sun.. This can be expected by the different algorithms used to retrieve the LoS velocity. SDO/HMI uses the so-called MDI-like method for its LoS observable, whereas in SO/PHI-HRT the LoS velocity field is a result of the RTE inversion. Different algorithms can also infer the physical quantities at slightly different heights depending on their response functions. Such a possibility was not considered in Schou et al. (2023), because the MDI-like algorithm was used for both instruments. The use of different algorithms in the SO/PHI-HRT and SDO/HMI pipelines for the retrieval of the velocities is possibly a further reason for the disagreement between the signals measured by these two instruments.
@@
As a last result, the Evershed flow within the sunspot penumbra is compared in Fig. 7. The velocity signal displayed on an elliptical polar grid is used to isolate the red- and blue-shifted parts of the penumbra. The comparison between SDO/HMI and SO/PHI-HRT, displayed in the right panel of Fig. 7, shows very good agreement between the two instruments within the penumbra boundary and a small divergence in the quiet Sun. The difference measured right outside the penumbra does not change significantly along the rotation angle axis, showing a small dependence over such a small FoV. Please note that the velocities plotted in blue have to be flipped in sign before this invariance becomes clear. This is in agreement with the result displayed in the bottom panel of Fig. 5, meaning that this difference is due to the general offset between the two instruments. Stated differently, the penumbra itself shows higher velocity and smaller difference with respect to the quiet Sun, which could be due to how MILOS and the MDI-like algorithm infer the velocity in different physical conditions. The smaller difference in the penumbra than in the quiet Sun can also explain the smaller spread in Fig. 6 compared to Fig. 3.

### v2

@@
The Polarimetric and Helioseismic Imager (SO/PHI) onboard Solar Orbiter is a spectropolarimeter scanning the Fe i line at 617.3 nm, providing data of the solar photosphere. The same line is sampled by the Helioseismic and Magnetic Imager (HMI) on board the Solar Dynamics Observatory (SDO) and many other on-ground instruments. In this paper, we aim at assessing the consistency between line-of-sight (LoS) velocity measurements from the two instruments. Reliable measurements of up and down flows from SO/PHI are crucial and unique when Solar Orbiter is facing the far side of the Sun. Also, a combination of measurements from two vantage points to study horizontal flows must rely on consistent observations. For this purpose, we compare the LoS velocity measured by SO/PHI’s High Resolution Telescope (SO/PHI-HRT) and SDO/HMI on 29 March 2023, when Solar Orbiter was crossing the Sun-Earth line at a distance of 0.39 au from the Sun. Because such co-alignments are rare, this configuration offered an almost unique opportunity to directly compare data products from both telescopes. The data are aligned and remapped to allow a pixel-by-pixel comparison of the whole time series of 4 hours length. Temporal and spatial variations are considered for a direct combination of the measurements. The LoS velocity distributions are evaluated and a clear linear relation is found between the two instruments with a slope of 0.96 and a correlation of 92%. We find that the signals form at similar heights, with a separation of 714 km, which is larger than previous estimates. A close-up look at the penumbra of a sunspot and its Evershed flow is presented. We conclude that the signals inferred by SO/PHI-HRT and SDO/HMI show very good agreement and high correlation when instrumental effects and large-scale velocities on the Sun are properly accounted for.
@@
Despite the clear advantages of satellite observations, such as the absence of the disturbance from the Earth’s atmosphere and the capability of uninterrupted data acquisition, many challenges have to be faced to achieve a consistent data reduction. In the case of SO/PHI the highly eccentric orbit and low telemetry periods when the spacecraft is at more than 1 au from Earth, make the SO/PHI data reduction considerably more challenging. On the other hand, SO/PHI can for the very first time measure the photospheric properties of the Sun from different viewpoints than Earth-bound observatories (Müller et al. 2020). This unique orbit opens the possibility for novel scientific applications, such as building fast synoptic maps of the LoS magnetic field (Loeschl et al. 2024b, a) or direct combination of intensity and magnetic field data in stereoscopic configuration (e.g. Valori et al. 2023; Albert et al. 2023; Romero Avila et al. 2024). Comparisons of the magnetic field vector between SO/PHI and SDO/HMI have been carried out by Sinjan et al. (2023) and Moreno Vacas et al. (2024) to assess the similarity between the data products of these two instruments. Such inter-instrument comparisons are fundamental for combining data from different vantage points, particularly when tracking features on the far side of the Sun as observed from Earth (e.g. Finley et al. 2025).
@@ 2 Data
The data products chosen for this comparison are the 45 s cadence data products of SDO/HMI (based on the so-called MDI-like algorithm, Couvidat et al. 2016) and the newly improved standard results of the SO/PHI-HRT pipeline (Sinjan et al. 2022; Kahil et al. 2022; Bailén et al. 2024), which is now also correcting for residual stray light, based on the inversion of the radiative transfer equations (RTE) in Milne-Eddington approximation with the MILOS code (Orozco Suárez & Del Toro Iniesta 2007). This data product became the standard data product in most recent data releases111Fifth SO/PHI-HRT data release notes. (see Sinjan et al. in preparation, for more details).
@@ 4.1 Fourier space
The power spectra of the remapped SO/PHI-HRT and SDO/HMI time series and their phase difference and coherence spectrum are shown in Fig. 2. The phase difference has been corrected for the time difference between the two time series. The coherence spectrum shows the linear relation between two signals and it is not affected by their amplitude (Vigeesh et al. 2017). The first thing to notice is the similarity between the power spectra of the two instruments, which is made more evident by the diagrams showing the phase difference and the coherence between the two time series. The phase difference remains consistently close to 0∘ across most of the domain, while the coherence spectrum shows very high values (above 80%) along the p-modes and f-mode ridges and in the internal gravity waves and granulation regime. The phase difference diagrams can be used to estimate the average height difference between the two signals (see equations (2) and (3) in Vesa et al. 2023). We select the area of the diagram where the waves have real vertical wavenumber, which means that they propagate and they are not evanescent (as when the vertical wave number is imaginary). We can then restrict our choice only to vertically propagating waves by considering only small horizontal wavenumbers (see the red dashed box in the bottom right panel in Fig. 2). We assume the cut-off frequency, buoyancy frequency, and sound speed to be equal to 5.4 mHz, 4.9 mHz, and 7 km/s respectively. We find that the SO/PHI-HRT signal forms approximately 7 km above SDO/HMI, with a broad distribution between -10 and 30 km. Previous result by Schou et al. (2023) suggested a formation height of ~156 km for SDO/HMI and ~158 km above the solar surface for SO/PHI obtained with the help of a MURaM simulation. This result was obtained by retrieving the velocity through the MDI-like algorithm also for SO/PHI, so we would expect those signals to be closer than those addressed in that work.
@@ 4.2 Distributions
As a second assessment of the similarity of the SDO/HMI and SO/PHI-HRT Dopplergrams, a scatter plot of the remapped SO/PHI-HRT and the non-interpolated SDO/HMI time series is shown in Fig. 3. Only the datasets with time difference of 5 s are selected for plotting in this figure because the width of the distribution is affected by the time discrepancy. The colors correspond to the number of points inside a specific bin in logarithmic scale. Two functions have been fit to this distribution: a linear function and a 2-D Gaussian distribution. The linear fit shows a clear correspondence between the data, with a slope of 0.9560.003, a correlation coefficient of 92%, and an offset of -282.91.2 m/s. The Gaussian distribution indicates the spread between the two signals, which is 116.61.0 m/s at 1. The uncertainties are estimated with the bootstrap method by repeating the fitting on 500 random sub-samples. We also determine the mean of the distribution, which provides information on the offset of the zero point between both time series. The average of the distribution for SDO/HMI is at -224 m/s, whereas the SO/PHI-HRT mean value is at -62 m/s. Discussing the absolute calibration of these instruments is a challenge that goes beyond the scope of this work, but it is indeed a point that needs to be mentioned. In the SDO/HMI case, the reference velocity can be set in many ways available in the literature (e.g. Mathew et al. 2003; Grigor’ev et al. 2011; Okamoto et al. 2009; de la Cruz Rodríguez et al. 2011; Welsch et al. 2013), including simply setting the mean signal at the center of the Sun to zero (after the removal of all the LSC). The same approach is not always possible for SO/PHI-HRT, due to its smaller FoV and the fact that it often points such that solar disk center does not lie in its FoV.
@@
To investigate the presence of any temporal dependence, we repeated the analysis described above for each of the 240 frames separately. Fig. 4 shows a selection of fit parameters as a function of time. The uncertainties are estimated with the bootstrap method by repeating the fitting on 80 random sub-samples. The top panel displays the results of the linear fit on each of the 240 observation pairs of frames. The blue points show the slope and the orange points the offset obtained by the linear fit. We can clearly distinguish a linear trend through the 4 hours of observation in which both values evolve towards a better concordance. The middle panel depicts the coordinates of the average of the Gaussian distribution. The left axis (in green, coordinates of the center of the distribution) corresponds to the mean values of each dataset of the remapped SO/PHI-HRT, the right coordinates (in red, coordinates of the center of the distribution) to SDO/HMI. The range of the two -axes is the same for both coordinates, meaning that the relative variation with time of both values is directly comparable. In this case, the SDO/HMI values do not show any significant variation with time, while the remapped SO/PHI-HRT values go from 70 to 55 m/s. The bottom panel shows the root mean square (rms) of the remapped SO/PHI-HRT (purple) and SDO/HMI (brown) time series, which is used to quantify the amplitudes measured by the instruments. The two axes in this case are exactly the same for both distributions. The remapped SO/PHI-HRT rms is overall higher than SDO/HMI as expected by the slope of the linear fit shown in Fig. 3 being smaller than one. On the other hand, we see a decrease of the rms in SO/PHI-HRT of ~5 m/s and an increase in SDO/HMI of ~10 m/s, which is also consistent with the increase of the slope in time.
@@
In a further step we look at the spatial variation of the results of the linear fit. For that, the FoV is divided into square regions of 88 pixels and we perform a linear fit in each of these regions through the whole time series. We also compute the averaged correlation between the two time series in each sub-region. The results are shown in Fig. 5. The top panel shows the spatial-distribution of the correlation, which is overall very high with a mean value of 93%. Lower values are found especially within the sunspot umbra, where the correlation drops to 70-80%. Many factors lead to this low correlation including the low intensity signal, the large line splitting due to the strong magnetic field that can become as large as the total wavelength range in the line that is scanned, the small velocity in the umbra, the few measured wavelength points within the line (which are different for SDO/HMI and SO/PHI), different algorithms to retrieve the velocity signal, all contribute to the low correlations between the two instruments within umbrae (see Sect. 5).
@@ 4.3 Sunspot penumbra
In this section the penumbra of the sunspot is analyzed in more detail. The penumbra is defined as those pixels with intensity higher than 55% and lower than 89% of the averaged intensity continuum in a nearby quiet Sun area. A reproduction of the scatter plot between the remapped SO/PHI-HRT and the SDO/HMI time series, focusing exclusively on the penumbra and on its Evershed flow (Evershed 1909) is visible in Fig. 6. The contours defining the penumbra boundaries are also visible in the inset. The uncertainties are estimated with the bootstrap method by repeating the fitting on 100 random sub-samples. It is evident that the scatter of the points is smaller than in Fig. 3, which can be explained by the smaller FoV considered here (and thus smaller spatial variation) and is well represented by the higher correlation coefficient (95%) with respect to that of the full FoV. Another remarkable result is the slope of the linear fit, which is comparable to the one obtained by fitting the data of the whole FoV within their uncertainties. Additionally, the offset between the two signals is smaller than that found in Fig. 3. This is expected from the position of the sunspot in the bottom panel of Fig. 5, where it lies in an area where the offset is smaller than the average (as obtained by the linear fit shown in Fig. 3).
@@
The middle panel of Fig. 7 displays the velocity on the elliptical polar grid. Note that the velocities plotted in this panel are temporal averages over the full 240 frames recorded by the SO/PHI-HRT. The two branches of the Evershed flow are clearly visible in the left and right part of the plot, which consists of the limb (outlined by the red box) and disk center (blue box) oriented sides of the penumbra, respectively. In order to compare the values obtained by the two instruments, we perform an azimuthal average (along the axis) in the red and blue boxes for each time series. The results are shown in the right panel of Fig. 7, where the axis represents the radial distance on the elliptical polar grid. The colors are the same as of the boxes shown in the middle panel, the crosses show the result obtained with the remapped SO/PHI-HRT and the circles those from SDO/HMI. The curves resulting from the data of the two instruments lie very close to each other until the penumbra-quiet Sun boundary (which lies roughly at a distance of 15 px along the axis). An offset, comparable to that shown in Fig. 6, is visible around the peaks of the velocity signals. The asymmetry between the two sides of the penumbra (limb- and disk center-side, e.g. Schlichenmaier et al. 2004) is to be expected. Also, a difference in the flows can be caused by the asymmetric shape of the sunspot.
@@ 5 Discussion
We have reported a comparison of the LoS velocity measured by SO/PHI-HRT and SDO/HMI while Solar Orbiter was in inferior conjunction. The data acquired by SO/PHI-HRT have been post-processed and remapped onto the SDO/HMI detector frame to allow a pixel-by-pixel comparison. Despite differences in instrumentation, calibration, cadence, and velocity retrieval methods, the comparison shows an excellent agreement of the LoS velocity signals obtained with these two telescopes. We measure a height difference of 714 km between the two time series and obtain a slope of 0.94 for the linear fit of the scatter plot between the two instruments and a correlation coefficient of above 90%. The measured height difference is larger than the result obtained from a MURaM simulation by Schou et al. (2023) and it is possible to estimate the velocity variation due to such a separation. If we consider the conservation of mass from one layer to the other, it holds
where is the density at each layer and is the vertical velocity. A ratio of 0.962 for 7 km height separation is obtained by considering the atmospheric densities from the VAL-C model (Vernazza et al. 1981), whereas 8 km of separation would be necessary to obtain a ratio equals to 0.956. Considering the broad distribution of the height separation between the two instruments, it is not possible to address it as the unique cause for having a ratio smaller than 1. One more reason is to be found in the bottom panel of Fig. 4, which shows that the rms of the remapped SO/PHI-HRT is always higher than that of SDO/HMI. This is due to the spatial deconvolution of the SO/PHI-HRT PSF including the stray light component during the data reduction, a correction that is not included in the SDO/HMI pipeline, but it can be made available on request. We compare these data with the velocities from SDO/HMI without stray light removed, as we want to compare the standard data products, which are the most likely to be used for studies comparing or combing velocities from the two instruments. One more element that could also contribute is the different sampling of the spectral line. SO/PHI-HRT samples the line in five points with a spectral resolution of 58000, SDO/HMI in six points with a spectral resolution of 81000. Stief et al. (2019) showed that the velocity retrieved with different sampling and at different points along the line can substantially affect the measurement. However, a lower spectral resolution reduces the variation in the retrieved velocity near the core of the line (see the yellow curve in Fig. 5 in Stief et al. 2019 for the SDO/HMI case), meaning that we do no expect this to be the main component of the difference between the two instruments.
@@
The spatial correlation and dependence of the linear fitting results are shown in Fig. 5. The correlation map shows very high values through the whole FoV with an average of 93%, with the exception of the sunspot umbra, where the measured LoS velocities are generally lower than in the penumbra or quiet sun. This can be expected by the different algorithms used to retrieve the LoS velocity. SDO/HMI uses the so-called MDI-like method for its LoS observable, whereas in SO/PHI-HRT the LoS velocity field is a result of the RTE inversion. Different algorithms can also infer the physical quantities at slightly different heights depending on their response functions. Such a possibility was not considered in Schou et al. (2023), because the MDI-like algorithm was used for both instruments. The use of different algorithms in the SO/PHI-HRT and SDO/HMI pipelines for the retrieval of the velocities is possibly a further reason for the disagreement between the signals measured by these two instruments.
@@
As a last result, the Evershed flow within the sunspot penumbra is compared in Fig. 7. The velocity signal displayed on an elliptical polar grid is used to isolate the red- and blue-shifted parts of the penumbra. The comparison between SDO/HMI and SO/PHI-HRT, displayed in the right panel of Fig. 7, shows very good agreement between the two instruments within the penumbra boundary and a small divergence in the quiet Sun comparable to the offset obtained with the linear fit shown in Fig. 6. The difference measured right outside the penumbra does not change significantly along the rotation angle axis, showing a small dependence over such a small FoV. Please note that the velocities plotted in blue have to be flipped in sign before this invariance becomes clear. This is in agreement with the result displayed in the bottom panel of Fig. 5, meaning that this difference is due to the general offset between the two instruments. Stated differently, the penumbra itself shows higher velocity and smaller difference with respect to the quiet Sun, which could be due to how MILOS and the MDI-like algorithm infer the velocity in different physical conditions. The smaller difference in the penumbra than in the quiet Sun can also explain the smaller spread in Fig. 6 compared to Fig. 3.